OpenStudy (anonymous):

I have 5 algebra math questions. If anyone helps i will fan and medal. :)

OpenStudy (anonymous):

Rewrite the radical as a rational exponent \[\sqrt[4]{7^{5}}\] A. \[7\frac{ 5 }{ 4 }\] B. \[7^{20}\] C. 7 D. \[7\frac{ 4 }{ 5 }\]

OpenStudy (anonymous):

@cutecaligirl_

OpenStudy (theopenstudyowl):

http://www.mathwords.com/r/radical_rules.htm

OpenStudy (anonymous):

i am going with D